OpenStudy (anonymous):

Wallace received a 28 percent discount on an item priced at $275. What is the total price that he paid for it? cost after discount = price × (1 – discount percentage)

OpenStudy (anonymous):

He paid $198 for the item.

OpenStudy (anonymous):

$275 /1 x 28/100 = 7700/100 = $77 is the amount of the discount. $275 - $77 =?